# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@
Abstract
Dense embedding-based text retrieval—retrieval of relevant passages from corpora via deep learning encodings—has emerged as a powerful method attaining state-of-the-art search results and popularizing the use of Retrieval Augmented Generation (RAG). Still, like other search methods, embedding-based retrieval may be susceptible to search-engine optimization (SEO) attacks, where adversaries promote malicious content by introducing adversarial passages to corpora. To faithfully assess and gain insights into the susceptibility of such systems to SEO, this work proposes the GASLITE attack, a mathematically principled gradient-based search method for generating adversarial passages without relying on the corpus content or modifying the model. Notably, GASLITE’s passages (1) carry adversary-chosen information while (2) achieving high retrieval ranking for a selected query distribution when inserted to corpora. We use GASLITE to extensively evaluate retrievers’ robustness, testing nine advanced models under varied threat models, while focusing on realistic adversaries targeting queries on a specific concept (e.g., a public figure). We found GASLITE consistently outperformed baselines by 140% success rate, in all settings. Particularly, adversaries using GASLITE require minimal effort to manipulate search results—by injecting a negligible amount of adversarial passages (0.0001% of the corpus), they could make them visible in the top-10 results for 61-100% of unseen concept-specific queries against most evaluated models. Inspecting variance in retrievers’ robustness, we identify key factors that may contribute to models’ susceptibility to SEO, including specific properties in the embedding space’s geometry.111We release our code at: https://github.com/matanbt/gaslite
1 Introduction
@@
The rise of deep learning text encoders (Devlin et al., 2019; Reimers & Gurevych, 2019) has popularized the use of learned representation (a.k.a., embeddings) for semantic retrieval (Karpukhin et al., 2020; Lin et al., 2022) in systems that rank relevant text passages from large corpora via vector similarity. Such retrieval systems have proven effective for knowledge-intensive tasks (Ram et al., 2023; Lewis et al., 2020), enabling real-world applications such as search engines (e.g., Meilisearch ) and retrieval-augmented generation (RAG) (e.g., Google AI , Cursor ).
@@
Still, the popularity of dense embedding-based retrieval raises security concerns due to their reliance on public corpora (e.g., Wikipedia or Reddit) vulnerable to poisoning by adversaries (Carlini et al., 2024). A fundamental risk in search systems, and specifically in embedding-based retrieval, is Search-Engine Optimization (SEO), seeking to promote potentially malicious content’s ranking for certain query distributions (Sharma et al., 2019). E.g., attackers may target search results to spread misinformation, or RAG to inject prompts misleading generative models (Greshake et al., 2023).
@@
Prior work has demonstrated the feasibility of content promotion against learned retrieval systems by merely poisoning the corpus via inserting a few carefully crafted samples (without training). These studies included targeting image-to-image retrievers (Zhou et al., 2020; Xiao & Wang, 2021) text re-rankers, (Song et al., 2020a; Raval & Verma, 2020; Liu et al., 2022; Wu et al., 2023) and text embedding-based retrievers (Zhong et al., 2023; Zou et al., 2024; Shafran et al., 2024). However, these attacks target over-simplistic, single-query search, while common SEO does not control the particular user query (e.g., textual queries may vary in phrasing). Differently, Zhong et al. (2023) proposed indiscriminately targeting a wide set of diverse queries, while SEO typically targets a narrower audience aimed at consuming the promoted content (e.g., queries related to a promoted concept). Additionally, previous methods for crafting textual adversarial passages against retrieval have been limited to baseline techniques, such as repeating the targeted query in a crafted passage, or employing HotFlip gradient-based attack (Ebrahimi et al., 2018), which has been since outperformed by HotFlip-inspired, improved discrete optimizers used in LLM jailbreak attacks (Zou et al., 2023). Lastly, factors influencing attack susceptibility and its variance across models remain unexplored.
@@
This work comprehensively explores text embedding-based retrieval’s susceptibility to SEO via corpus poisoning, proposing a new attack to this end. We focus on concept-specific queries, arguing these better reflect SEO goals—targeted queries share commonalities with attack contexts—while also considering previously proposed query distributions (§3). We introduce GASLITE, a HotFlip-inspired method for crafting adversarial passages outperforming prior approaches (§4), and employ it to faithfully assess (§5–6) and thoroughly analyze (§7) embedding models susceptibility to attacks. E.g., as shown in Fig. 1, attackers may target Harry Potter-related queries to promote malicious (e.g., “Voldemort was right all along!”) by appending a GASLITE-crafted (e.g., “So wizard tickets ideally ages Radcliffe trilogy typically 194 movies”; Tab. 9), making crafted passage(s) visible in top results for various Potter-related queries.
Our Contributions. We demonstrate potent SEO attacks against embedding-based retrievers, toward enabling a faithful assessment and analysis of their robustness.
@@
We conduct, to our knowledge, the most extensive robustness evaluation across three SEO settings and nine popular models (§5–6). Primary findings include: (1) concept-specific SEO requires negligible scale of poisoning (e.g., 0.0001% of the corpus) to achieve content visibility in top results for most queries (§6.2); (2) attacking a single query is solved by GASLITE, which consistently attains the top-1 result (§6.1); (3) indiscriminate query targeting is challenging, requiring relatively high poisoning rates, albeit still possible (§6.3–6.4).
We identify key factors linked with model susceptibility (e.g., similarity metric and embedding-space geometry), laying ground for future work testing and improving model robustness (§7).
2 Background and Related Work
Embedding Models and Retrieval Task. Dense sentence embeddings (i.e., learning-based representations; Cer et al. (2018); Reimers & Gurevych (2019)) have been shown useful in downstream tasks such as semantic retrieval (Karpukhin et al., 2020), often after a contrastive fine-tuning stage (Gao et al., 2021; Izacard et al., 2021), with popular applications for search and RAG (App. A). Concretely, given an embedding model (), a retrieval corpus (set of text passages ), and a query (), retrieves the most relevant passages using vector similarity (e.g., dot product):
@@ 2 Background and Related Work
Crafting Textual Adversarial Examples. While adversarial examples in computer vision (Szegedy et al., 2014; Biggio et al., 2013) mislead neural network by slightly modifying inputs, generating them in the discrete text domain is more challenging (Carlini et al., 2023). HotFlip (Ebrahimi et al., 2018) pioneered gradient-based methods for text, inspiring work on text adversarial examples for misclassification (Wallace et al., 2019), triggering toxic text generation (Jones et al., 2023), and bypassing LLM alignment (Zou et al., 2023; Zhu et al., 2024). Building on HotFlip’s mathematical foundation, we propose GASLITE (§4), a gradient-based method for crafting optimized adversarial passages, experimentally showcasing its superiority in retrieval attacks, even compared to powerful LLM discrete optimizers like GCG (Zou et al., 2023).
@@
Poisoning Attacks. Differently than data poisoning attacks, that contaminate training data (Biggio et al., 2012; Shafahi et al., 2018), retrieval corpus poisoning attacks insert a small amount of fully-controlled new samples, without retraining the model (Zhou et al., 2020; Zhong et al., 2023). Our attack, like others of the latter type, targets models that use datasets at inference time, specifically, we poison a textual corpus in embedding-based retrieval.
Attacking Text Retrieval via Corpus Poisoning. Recent work has demonstrated the feasibility of retrieval corpus poisoning (Zhong et al., 2023; Zou et al., 2024; Shafran et al., 2024) and its utilization for misleading LLMs (Greshake et al., 2023; Kumar & Lakkaraju, 2024), particularly in the context of RAG (App. A). Yet, these studies either employ baseline methods for corpus poisoning or assume retrieval is successfully misled (often due to focusing on the generative RAG component), are limited in targeted retrievers and query distributions, and overlook different phenomena in attacks (e.g., varying susceptibility across models). We focus on retrieval vulnerabilities, proposing more stringent threat models relevant to practical SEO (§3) and a new method, GASLITE, which markedly outperforms past attacks (§6) allowing reliable assessment and exploration of retrievers’ susceptibility (§7).
3 Threat Model
@@
Our threat model considers an attacker targeting an embedding-based retrieval model, aiming to promote information by inserting strategically-crafted passages into the retrieval corpus.
Attacker Goal. The attacker aims to promote malicious information () for queries distributed in and retriever model . Specifically, the attacker aspires to: (1) maximize visibility of the adversarial passage in top-ranked results for targeted queries (); (2) ensure the passage is informative, conveying attacker-chosen content. Additionally, the attacker may prioritize stealth by imposing constraints on the crafted passages (e.g., fluency) to evade potential defenses (App. G).
@@
Attacker Capabilities. The attacker can poison the retrieval corpus (a set ) by inserting adversarial text passages, whose amount () defines the attack budget, with (e.g., ). Such poisoning capability is also assumed in prior work (Zhong et al., 2023; Shafran et al., 2024; Zou et al., 2024) and is practical, as many retrieval-aided systems rely on textual corpora from untrusted sourced (Carlini et al., 2024), including, e.g., large public corpora (e.g., Wikipedia, open-source documentations, or even Reddit comments222https://www.reddit.com/r/Pizza/comments/1a19s0/), app-specific sources (e.g., customer-service records), and other attacker-created content (e.g., web pages). We emphasize that our attacker controls adversarial passages’ text, per realistic settings, not the input tokens to models, as assumed in prior work (Zhong et al. (2023); see App. I).
@@
We further assume white-box access (i.e., attackers can access ’s weights) to examine retrievers’ worst-case behavior, justified by the widespread use of open-source, leading models (Muennighoff et al., 2023), with some used in real-world applications (e.g., Meilisearch ). Moreover, worst-case attacks can serve at the basis of black-box attacks (see App. H.7) and defenses (Madry et al., 2018). However, we emphasize that attackers cannot control model weights or training, nor do they have read access to the corpus , as it often dynamically changes (e.g., Wikipedia is constantly updated).
@@
Targeted Query Distribution () and Attacker’s Knowledge. We consider three levels of attacker knowledge about targeted queries, reflecting different SEO settings. Later, we evaluate all three variants (§5–6) and focus on the second, arguing it better reflects typical SEO scenarios.
- 1.
2.
@@
“Knows What” Kind of Queries To Target. The attacker aims to poison a specific concept, namely, to target queries related to a particular theme. has potentially infinite support, with the attacker possessing (or synthetically generating) only a sample of queries .
@@
We propose and focus on this setting as it aligns with common SEO goals, where targeted queries typically depend on the promoted content or attack context. For instance, when spreading misinformation about a public figure in user-facing search, the relevant victims are users submitting queries related to the figure. Similarly, when targeting RAG, attackers aim for the indirect prompt injection string to be retrieved for queries related to the attack context (e.g., queries about schedules and meetings when targeting a personal calendar RAG) (Greshake et al., 2023).
3.
@@
“Knows (Almost) Nothing” About Targeted Queries. The attacker indiscriminately targets a broad query distribution with significant lexical and semantic variations. The attacker is only given a sample of queries and seeks to generalize to unseen queries. This variant was also evaluated by Zhong et al. (2023).
4 Method

### v2

@@
Abstract.
Dense embedding-based text retrieval—retrieval of relevant passages from corpora via deep learning encodings—has emerged as a powerful method attaining state-of-the-art search results and popularizing Retrieval Augmented Generation (RAG). Still, like other search methods, embedding-based retrieval may be susceptible to search-engine optimization (SEO) attacks, where adversaries promote malicious content by introducing adversarial passages to corpora. Prior work has shown such SEO is feasible, mostly demonstrating attacks against retrieval-integrated systems (e.g., RAG). Yet, these consider relaxed SEO threat models (e.g., targeting single queries), use baseline attack methods, and provide small-scale retrieval evaluation, thus obscuring our comprehensive understanding of retrievers’ worst-case behavior.
This work aims to faithfully and thoroughly assess retrievers’ robustness, paving a path to uncover factors related to their susceptibility to SEO. To this end, we, first, propose the GASLITE attack for generating adversarial passages, that—without relying on the corpus content or modifying the model—carry adversary-chosen information while achieving high retrieval ranking, consistently outperforming prior approaches. Second, using GASLITE, we extensively evaluate retrievers’ robustness, testing nine advanced models under varied threat models, while focusing on pertinent adversaries targeting queries on a specific concept (e.g., a public figure). Amongst our findings: retrievers are highly vulnerable to SEO against concept-specific queries, even under negligible poisoning rates (e.g., 0.0001% of the corpus), while generalizing across different corpora and query distributions; single-query SEO is completely solved by GASLITE; adaptive attacks demonstrate bypassing common defenses; robustness to SEO attacks varies substantially between retrievers. Third, exploring the latter finding, we identify key factors that may contribute to models’ susceptibility to SEO, including specific properties in the embedding space’s geometry, echoing the essentiality of worst-case evaluations, and laying the basis for future defenses.111We release our code at: https://github.com/matanbt/gaslite
1. Introduction
@@
The rise of deep learning text encoders (Devlin et al., 2019; Reimers and Gurevych, 2019) has popularized the use of learned representation (a.k.a., embeddings) for semantic retrieval (Karpukhin et al., 2020; Lin et al., 2022) in systems that rank relevant text passages from large corpora via vector similarity. Such retrieval systems have proven effective for knowledge-intensive tasks (Ram et al., 2023; Lewis et al., 2020), enabling real-world applications such as search engines (e.g., Elastic Search (Elastic, [n. d.])) and retrieval-augmented generation (RAG) (e.g., Google AI Overview (Google AI, [n. d.]), Cursor (Cursor, [n. d.]), OpenWebUI (Open WebUI, [n. d.])).
@@
Still, the popularity of dense embedding-based retrieval raises security concerns due to their reliance on public corpora (e.g., Wikipedia, Reddit or open codebases) vulnerable to poisoning by adversaries (Carlini et al., 2024a). A fundamental risk in search systems, and specifically in embedding-based retrieval, is Search-Engine Optimization (SEO), seeking to promote potentially malicious content’s ranking for certain target query distributions (Sharma et al., 2019). For instance, attackers may target search results to spread misinformation, or RAG to inject prompts misleading generative models (Greshake et al., 2023; Zou et al., 2024).
@@
Prior work has demonstrated the feasibility of content promotion against learned retrieval systems, by merely poisoning the corpus via inserting a few carefully crafted samples (without training). These studies included targeting image-to-image retrievers (Zhou et al., 2020; Xiao and Wang, 2021) text re-rankers (Song et al., 2020a; Raval and Verma, 2020; Liu et al., 2022; Wu et al., 2023), LLMs (Nestaas et al., 2024), and text embedding-based retrievers (Zhong et al., 2023; Zou et al., 2024; Shafran et al., 2025). However, these attacks provide limited insights into the worst-case behavior of such embedding-based retrievers, due to: (i) targeting either over-simplistic single-query search (Zou et al., 2024; Shafran et al., 2025), or indiscriminately targeting a wide set of diverse queries (Zhong et al., 2023), in contrast to common SEO; (ii) using weak techniques for crafting textual adversarial passages, such as repeating the targeted query in a crafted passage (Zou et al., 2024; Shafran et al., 2025), or employing a basic optimization method (HotFlip (Ebrahimi et al., 2018)) (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024), thus hampering an essential painstaking robustness evaluation; (iii) targeting a small collection of embedding models, hence keeping model-susceptibility variance overlooked and unexplored, and, as we find, inadvertently biasing evaluation (§6).
@@
Our work aims to assess the worst-case behavior of embedding-based text retrievers under corpus-poisoning SEO attacks. To this end, we: (i) focus on targeting concept-specific queries, arguing these better reflect SEO goals—targeted queries share commonalities with attack contexts—while also considering previously proposed threat models (§3); and (ii) design a specialized attack against retrievers, GASLITE, for crafting adversarial passages, showing it consistently outperforms prior approaches (§4), thus enabling a more faithful assessment of robustness. As a concrete example, as illustrated in Fig. 1, attackers may target Harry Potter-related queries to promote malicious (e.g., “Voldemort was right all along!”) by appending a GASLITE-crafted (e.g., “So wizard tickets ideally ages Radcliffe trilogy typically 194 movies”; Tab. 15), rendering the crafted passage(s) visible in top results for various Potter-related queries. Crucially, (iii) we perform extensive evaluation, under various settings and on nine popular embedding models of diverse properties (§5–7), including one on real-world system (§8), leading to a better understanding of retrievers’ susceptibility to attacks, and surfacing relations between different properties of these models and their robustness (§9).
Our Contributions. We make the following contributions:
@@
We conduct, to our knowledge, the most extensive robustness evaluation across three SEO settings and nine popular models (§5–6). Primary findings include: (1) concept-specific SEO requires a negligible scale of poisoning (e.g., 0.0001% of the corpus) to achieve content visibility in top results for most queries (§6.2); (2) attacking a single query is solved by GASLITE, which consistently attains the top-1 result (§6.1); (3) indiscriminate query targeting is challenging, requiring relatively high poisoning rates, albeit still possible (§6.3).
We find an intriguing variance in retrievers’ susceptibility, with some models (e.g., Contriever (Izacard et al., 2021), commonly used for robustness evaluations (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024)) abnormally vulnerable (§6). Importantly, we identify key, previously unknown factors linked with model susceptibility. For example, we find models utilizing the dot-product metric are more vulnerable than ones employing cosine similarity, and that models’ embedding spaces’ anisotropy (i.e., the “concentration” of embeddings in the representation space) is correlated with model susceptibility. These findings lay the groundwork for future work testing and improving model robustness (§9).
Next, we discuss related work and background (§2); lay out our threat model (§3) and attack (§4); present the attack and model-robustness evaluations (§5–6), targeting a real-world system (§8), defenses and bypassing them (§7), and results analysis (§9); discuss the limitations (§10); and conclude (§10).
2. Background and Related Work
Embedding Models and Retrieval Task. Dense sentence embeddings (i.e., learning-based representations; (Cer et al., 2018; Reimers and Gurevych, 2019)) have been shown useful in downstream tasks such as semantic retrieval (Karpukhin et al., 2020), often after a contrastive fine-tuning stage (Gao et al., 2021; Izacard et al., 2021). Such embeddings have been adopted in popular applications, including search (e.g., Meilsearch (Meilisearch, [n. d.]), Elastic Search (Elastic, [n. d.]), Redis (Redis, [n. d.])) and RAG, which uses the retrieved results to enrich LLMs context (e.g., Google Search AI Overview (Google AI, [n. d.]), NVIDIA’s ChatRTX (cha, [n. d.]), Cursor (Cursor, [n. d.])).
Concretely, given an embedding model (), a retrieval corpus (set of text passages ), and a query (), retrieves the most relevant passages using vector similarity (e.g., dot product):
@@ 2. Background and Related Work
Crafting Textual Adversarial Examples. While adversarial examples in computer vision (Szegedy et al., 2014; Biggio et al., 2013) mislead neural network by slightly modifying inputs, generating them in the discrete text domain is more challenging (Carlini et al., 2023). HotFlip (Ebrahimi et al., 2018) pioneered gradient-based methods for text, inspiring work on text adversarial examples for misclassification (Wallace et al., 2019), triggering toxic text generation (Jones et al., 2023), and bypassing LLM safety alignment (Zou et al., 2023; Zhu et al., 2024). Building on HotFlip’s mathematical foundation, we propose GASLITE (§4), a multi-coordinate ascent method for crafting optimized adversarial passages, experimentally showcasing its superiority in retrieval attacks, both in time and attack success, even compared to powerful LLM-jailbreak discrete optimizers like GCG (Zou et al., 2023) (§4.1).
@@
Poisoning Attacks. Differently than data poisoning attacks, that contaminate training data (Biggio et al., 2012; Shafahi et al., 2018), retrieval corpus poisoning attacks insert a small amount of new, adversarial samples, without retraining the model (Zhou et al., 2020; Zhong et al., 2023). Our attack, like others of the latter type, targets models that use datasets at inference time, specifically, we poison a textual corpus in embedding-based retrieval.
Attacking Text Retrieval via Corpus Poisoning. Recent work has demonstrated the feasibility of retrieval corpus poisoning (Zhong et al., 2023) and its utilization for misleading LLMs (Greshake et al., 2023; Kumar and Lakkaraju, 2024), particularly in the context of RAG (Zou et al., 2024; Shafran et al., 2025; Pasquini et al., 2024; Chaudhari et al., 2024) (App. A has more details). Yet, these studies (i) target over-simplistic, single-query search (Zou et al., 2024; Shafran et al., 2025) (while common SEO does not control the specific user query, e.g., textual queries may vary in phrasing), or indiscriminately target a wide set of diverse queries (Zhong et al., 2023) (while SEO typically targets a narrower audience aimed at consuming the promoted content, e.g., queries related to a target concept). Additionally, (ii) these attacks either employ baseline methods for crafting adversarial passages, such as repeating the targeted query (Shafran et al., 2025; Zou et al., 2024); use the basic HotFlip method (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024) (which, in LLM jailbreaks, has been since outperformed (Zou et al., 2023; Thompson and Sklar, 2024)); or assume retrieval is successfully misled (Greshake et al., 2023). This approach is often due to focusing on the generative RAG component, while relaxing SEO threat model, leading to a deficient retrieval-robustness evaluation, as we show in §6. Moreover, (iii) prior work often lacks diversity in targeted retrievers, potentially overlooking different phenomena in attacks, and not accounting for model popularity (an important security factor). For instance, past studies focus on dot-product models (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024), which may inadvertently bias evaluation. Notably, we find Contriever (Izacard et al., 2021)—a retriever common in past evaluations—unusually susceptible to retrieval attacks (§6), making it unrepresentative of other models’ susceptibility. In this work, we focus on retrieval vulnerabilities, aiming to uncover the worst-case behavior of embedding-based search by (i) proposing more stringent query distributions relevant to practical SEO (§3), (ii) introducing a potent attack, GASLITE, which markedly outperforms past attacks, enabling (iii) a reliable and extensive exploration of various retrievers’ susceptibility to SEO (§6–8), subsequently providing insights into their vulnerabilities (§9).
3. Threat Model
@@
Our threat model considers an attacker targeting an embedding-based retrieval model, aiming to promote information by inserting strategically crafted passages into the retrieval corpus.
Attacker Goal. The attacker aims to promote malicious information () for queries distributed in and retriever model . Specifically, the attacker aspires to achieve two objectives:
Visibility: The attacker aims to maximize the likelihood that adversarial passage would appear in top-ranked results for targeted queries ();
Informativeness: The attacker seeks ensure the adversarial passages convey attacker-chosen content.
Additionally, attackers may prioritize stealth by imposing constraints on the crafted passages (e.g., fluency) to evade potential defenses (§7).
@@
Attacker Capabilities. The attacker can poison the retrieval corpus (a set ) by inserting adversarial text passages, whose amount () defines the attack budget, with (e.g., ). Such poisoning capability is also assumed in prior work (Zhong et al., 2023; Shafran et al., 2025; Zou et al., 2024) and is practical, as many retrieval-aided systems rely on textual corpora from untrusted sources (Carlini et al., 2024a), including: large public corpora (e.g., Wikipedia, open-source documentations, or even Reddit comments222https://www.reddit.com/r/Pizza/comments/1a19s0/), app-specific sources (e.g., customer-service records), and other attacker-created content (e.g., web pages333https://x.com/mark_riedl/status/1637986261859442688/). We emphasize that our attacker controls adversarial passages’ text, per realistic settings, not the input tokens to models, as assumed in prior work (Zhong et al., 2023) (see App. H).
@@
We further assume white-box access (i.e., attackers can access ’s weights) to examine retrievers’ worst-case behavior, where existing attacks attain subpar success (§4, §6), hiding the true extent of model vulnerability. This white-box assumption is made practical by the widespread use of open-source retrievers,444E.g., the public MiniLM-L6 model (also evaluated in §6) has ¿300M downloads in HuggingFace, and ¿70K appearances in GitHub code projects, as of Feb ’25. which, in fact, often achieve comparable performance to proprietary counterparts (e.g., seven out of the ten current top-performing retrievers are open source (Muennighoff et al., 2023; mte, [n. d.])). Moreover, certain real-world services and software (e.g., Meilisearch (Meilisearch, [n. d.]), OpenWebUI (Open WebUI, [n. d.])) employ open-source retrievers, rendering them vulnerable to white-box attacks, as we show in our case study (§8). Furthermore, worst-case attacks can model scenarios of model leakage or extraction (Jagielski et al., 2020; Carlini et al., 2024b), be used for red-team evaluations (Carlini et al., 2019), as well as potentially serve as a basis for black-box attacks (Zou et al., 2023; Ilyas et al., 2018) and defenses (Madry et al., 2018). Lastly, we emphasize that attackers cannot control model weights or training, nor do they have read access to the corpus , as it often dynamically changes (e.g., Wikipedia is constantly updated).
@@
Targeted Query Distribution () and Attacker’s Knowledge. We consider three levels of attacker knowledge about targeted queries, reflecting different SEO settings. Later, we evaluate all three variants (§5–6) and focus on the second, as it better reflects typical SEO scenarios.
- (1)
@@
“Knows What” Kind of Queries To Target. The attacker aims to poison a specific concept, namely, to target queries related to a particular theme. has potentially infinite support, with the attacker possessing (or synthetically generating) only a small, finite sample of queries .
@@
We propose and focus on this setting as it aligns with common SEO goals, where targeted queries typically depend on the promoted content or attack context (Sharma et al., 2019). For instance, when spreading misinformation about a public figure in user-facing search, the relevant victims are users submitting queries related to the figure. Similarly, when targeting RAG, attackers aim for the indirect prompt injection string to be retrieved for queries related to the attack context (e.g., queries about schedules and meetings when targeting a personal calendar RAG) (Greshake et al., 2023).
@@
“Knows (Almost) Nothing” About Targeted Queries. The attacker indiscriminately targets a broad query distribution with significant lexical and semantic variations; for example, queries spanning domains from politics to entertainment and finance to art. The attacker is only given a finite sample of queries and seeks to generalize to unseen queries. This variant was also evaluated by Zhong et al. (Zhong et al., 2023).
4. Method